OpenStudy (anonymous):

2a + 3b - a simplify. show your work.

OpenStudy (anonymous):

a + 3b

OpenStudy (anonymous):

(2a-a)+3b =a +3b

OpenStudy (aravindg):

(2a-a)+3b =a +3b